A gaseous oxide contains 30.4% of nitrogen, 1 molecule of which contains 1 N atom. The density of the oxide relative to oxygen is ?

Grade:12

1 Answers

Sunil Kumar FP
10 years ago
the formula of the compound will be NO2.
let tye foemula be NOx
% of nitrogen =14/14+16x=.304
x=2
No2
ratio ofvtheir realtive density is directly proportional to their molecular mass
theirfore ratio =46/32=23/16

tyanks and regards
sunil kr
askIItian faculty